OpenStudy (anonymous):

Expand and simplify the following expression. If possible, write your answer without exponents. log4(.25(x+7)^6)

OpenStudy (unklerhaukus):

\[ \log_4(.25(x+7)^6)=\log_40.25+\log_4(x+7)^6\] \[=\log_4\frac 14+6\log_4(x+7)\] \[=\log_44^{-1}+6\log(x+7)\] \[=-1+6\log(x+7)\]

OpenStudy (anonymous):

There are so many rules for logs and my brain rejects them. Thank you.